# Superpixel Transformers for Efficient Semantic Segmentation

### v1

@@ Abstract
Semantic segmentation, which aims to classify every pixel in an image, is a key task in machine perception, with many applications across robotics and autonomous driving. Due to the high dimensionality of this task, most existing approaches use local operations, such as convolutions, to generate per-pixel features. However, these methods are typically unable to effectively leverage global context information due to the high computational costs of operating on a dense image. In this work, we propose a solution to this issue by leveraging the idea of superpixels, an over-segmentation of the image, and applying them with a modern transformer framework. In particular, our model learns to decompose the pixel space into a spatially low dimensional superpixel space via a series of local cross-attentions. We then apply multi-head self-attention to the superpixels to enrich the superpixel features with global context. Finally, we directly project the superpixel features back into the pixel space using the associations between the superpixels and the image pixel features. Reasoning in the superpixel space allows our method to be substantially more computationally efficient compared to convolution-based decoder methods. Yet, our method achieves state-of-the-art performance in semantic segmentation due to the rich superpixel features generated by the global self-attention mechanism. Our experiments on Cityscapes and ADE20K demonstrate that our method matches the state of the art in terms of accuracy, while outperforming in terms of model parameters and latency.
@@ I Introduction
Figure 1: Our method enables efficient segmentation of high-resolution camera images by learning to decompose the images into a set of superpixels. Specifically, we oversegment the image pixels into a small set of soft superpixels (left) via a series of local cross attentions. The superpixels are then refined via a set of multi-head self attentions. Finally, we fuse the features with the superpixel-pixel associations to produce a dense semantic segmentation (right).
Figure 2: Our proposed Superpixel Transformer architecture. Given an image, we first generate hypercolumn features with an off-the-shelf encoder backbone. Our superpixel tokenization module uses a series of local dual-path cross-attentions to generate features for each superpixel. The superpixel features are then enriched by several multi-head self-attention (MHSA) layers, while the associations between each superpixel and pixel feature are computed from their respective features. Finally, the superpixel predictions are unfolded into the dense pixel space using the associations. Note that the figure illustrates a hard assignment between pixels and superpixels for simplicity, while in practice we apply a differentiable soft assignment.
@@
Operating on the superpixel level provides a number of notable benefits. Conventional pixel-based approaches are limited by the high dimensionality of the pixel space, making global self-attention computationally intractable. Numerous approaches such as axial attention [42] or window attention [31] have been developed to work around these issues by relaxing the global attention to a local one. By over-segmenting the image into a small set of superpixels, we are able to efficiently apply global self-attention on the superpixels, providing full global context to the superpixel features, even when reasoning about high-resolution images. Despite applying global self-attention (vs conventional convolutional neural networks) in our model, our method is more efficient than existing methods due to the low dimensionality of the superpixel space. Finally, we are able to directly project the superpixel features back to the pixel space, using the associations between the superpixel and pixel features.
@@
  • A novel network architecture that uses local cross-attention to significantly reduce the spatial dimensionality of pixel features to a small set of superpixel features, enabling learning the global context between them.
  • A superpixel association and unfolding scheme that projects each superpixel feature back to a dense pixel segmentation, discarding the CNN pixel decoder.
@@ II Related Work
Transformers are also used as the decoders [2] for image segmentation. A popular design is to generate masks embedding vectors from object queries and then multiply them with the pixel features to generate masks [39, 44]. For example, MaX-DeepLab [41] proposes an end-to-end mask transformer framework that directly predicts class-labeled object masks. Segmenter [38] and MaskFormer [12] tackle semantic segmentation from the view of mask classification. K-Net [52] generates segmentation masks by a group of learnable kernels. Inspired by the similarity between mask transformers and clustering algorithms [33], clustering-based mask transformers are proposed to segment images [49, 48, 50]. Deformable transformer [57] is also used for improving the image segmentation as in Panoptic SegFormer [30] and Mask2Former [11]. Region Proxy [53] also incorporates the idea of superpixels into a deep segmentation network by using a CNN decoder to learn the association between each pixel and superpixel.
All of the prior works that apply transformers to segmentation have, in some way, relied on a dense CNN decoder to generate the final dense features, and then combined these features with an attention mechanism to improve performance. Our method, in comparison, uses cross attention to reduce the image into a small set of superpixels, and only applies self-attention in this superpixel space in the decoder. This allows our method to operate on a significantly lower dimensional space (often 32x smaller than the image resolution), while utilizing the benefits that come with global self-attention to achieve state-of-the-art performance.
@@ III Method
    Superpixel Self-Attention: A series of multi-head self-attention layers to refine the superpixel features.
@@
    Superpixel Association: Associating the refined superpixel and pixel features to obtain the final dense semantic segmentation.
@@ III-A Pixel Feature Extraction
Typical convolutional neural networks, such as ResNet [21] and ConvNeXt [32], are employed as the encoder backbone. On top of the encoder output, we apply a multi-layer perceptron (MLP), and bilinear resize to the features after stage-1 (stride 2), stage-3 (stride 8), and stage-5 (stride 32). The multi-scale features are combined with addition to form hypercolumn features [20]. Each pixel feature is then represented by their corresponding hypercolumn features, which are fed to the following Superpixel Tokenization module.
@@ III-B Superpixel Tokenization
where $\mathcal{N}(x)$ denotes the neighborhood of $x$. For each superpixel $S_{i}$, there are $9\cdot h\cdot w$ pixel neighbors, where $[h,w]$ is the size of the patch covered by one superpixel, while each pixel has 9 superpixel neighbors. We illustrate this neighborhood in Figure 3. The local dual-path cross-attention repeats $n$ times to generate the output superpixel, $S_{i}^{t_{n}}$ and pixel, $I_{p}^{t_{n}}$, features.
We note that our proposed local dual-path cross-attention aims to serve three purposes:
@@
III-C Superpixel Self-Attention
Given the updated superpixel features from the Superpixel Tokenization module, we apply $k$ multi-head self-attention (MHSA) layers [40] to learn global context information between superpixels, producing outputs $F_{i}$. Performing MHSA on the superpixel features is significantly more efficient than on the pixel features, since the number of superpixels is much smaller. In our experiments, we typically use a superpixel resolution that is 32x smaller than the input resolution. Finally, as opposed to the CNN pixel decoder used in other approaches [10, 12], our enriched superpixel features $F_{i}$ can be directly projected back to the final pixel-level semantic segmentation output without any additional layers, as described in Section III-D.
@@ III-D Superpixel Association
To project the superpixel features back into the pixel space, we use the outputs of the Superpixel Tokenization module, $I_{p}^{t_{n}}$ and $S_{i}^{t_{n}}$, to compute the association between each pixel and its 9 neighboring superpixels:
@@
The final dense semantic segmentation, $Y$, is then computed at each pixel, $p$, as the combination of each refined superpixel feature output from the Superpixel Self-Attention module, $F_{i}$, weighted by the above associations:
$Y_{p}=$ $\sum_{i\in\mathcal{N}(p)}Q_{pi}\cdot F_{i}.$ (6)
@@
MaskFormer [12] ResNet-50 [21] FPN [lin2017feature] - - - 78.5
Mask2Former[11] ResNet-50 [21] MSDeformAttn [57] - - - 79.4
Panoptic-DeepLab [10] ResNet-50 [21] DeepLabv3+ [8] 43M 517G - 78.7
$*{}^{*}$ [53] ViT-S [17] RegProxy [53] 23M 270G - 79.8
SP-Transformer ResNet-50 [21] Superpixel-Association 29M 253G 15.3 80.4
$k$MaX-DeepLab [50] ResNet-50 [21] Axial-Decoder [42] 56M 434G 9.0 79.7
MaskFormer [12] ResNet-101c - - - - 81.4
${}^{\dagger}$ [11] Swin-L [31] MSDeformAttn [57] - - - 83.3
$*{}^{*}$ [53] ViT-L/16 [17] RegProxy [53] 307M - - 81.4
SegFormer [47] MiT-B5 [47] All-MLP-Decoder [47] 85M 1,448G 2.5 82.4
SP-Transformer ConvNeXt-L [32] Superpixel-Association 202M 1,557G 3.6 83.1
$k$MaX-DeepLab [50] ConvNeXt-L [32] Axial-Decoder [42] 232M 1,673G 3.1 83.5
@@
Method Backbone Pixel decoder Crop Params $\downarrow$ FLOPs $\downarrow$ FPS $\uparrow$ mIoU $\uparrow$
RegProxy [53] ViT-Ti/16 [17] RegProxy [53] 512 6M 3.9G 38.9 42.1
MaskFormer [12] ResNet-50 [21] FPN [lin2017feature] 512 41M 53G 24.5 44.5
SP-Transformer ResNet-50 [21] Superpixel-Association 640 29M 78G 40.8 43.7
$k$MaX-DeepLab [50] ResNet-50 [21] Axial-Decoder [42] 641 57M 75G 38.7 45.0
RegProxy [53] ViT-L/16 [17] RegProxy [53] 640 306M 326G 6.6 52.9
[…]
SP-Transformer ConvNeXt-L [32] Superpixel-Association 640 202M 325G 14.9 53.0
$k$MaX-DeepLab [50] ConvNeXt-L [32] Axial-Decoder [42] 641 232M  [^3] G 14.0 54.8
TABLE II: ADE20K val set results. We evaluate FLOPs and FPS with input $640\times 640$ for SP-Transformer on a Tesla V100-SXM2 GPU. Our method outperforms the prior RegProxy work, while remaining competitive with other prior works, and at the highest FPS. $\dagger$ Pretrained on ImageNet-22k.
@@ IV-A Implementation Details
Superpixel Self-Attention 4 multi-head attention layers are applied in the superpixel self-attention stage, with 4 heads each outputting 256 channels in each layer. A final MLP is applied to project the refined superpixel features into semantic predictions (i.e., the output channel is equal to the number of classes in the target dataset).
@@ IV-B1 Cityscapes Dataset
Table I compares the results of our Superpixel Transformer model to other transformer-based state-of-the-art models for semantic segmentation on the Cityscapes val set. In addition to other semantic segmentation models, we also compare against the state of the art panoptic segmentation method, kMaX-DeepLab [50]. While this method is trained on a slightly different task, we find that, for computational cost, it is the most fair comparison, as our training schedule and pipeline is most similar to theirs.
The strength of our method is most evident for the smaller ResNet-50 backbone, where our model is roughly half the size and latency of kMaX-DeepLab, while improving upon mIoU by 0.6 against the previous SOTA, RegProxy [53]. As the ResNet-50 backbone is relatively small, most existing models are dominated by the size of their decoders, which allows our model’s reduction in decoder size to have significant impacts on the overall size and performance, with a $70\%$ improvement in FPS compared to kMaX-DeepLab. In addition, we expect to see this effect even more for even smaller backbones.
For the larger ConvNeXt-L backbone, our model has a similar absolute reduction in params and FLOPs, as compared to the equivalent kMaX-DeepLab model. However, as the model is largely dominated by the size of the backbone, the overall improvements are more modest. Nonetheless, our model is able to achieve near state-of-the-art performance, especially compared to the prior semantic segmentation methods, where we outperform most of the prior works, except for Mask2Former, where we are within 0.2mIoU. We note that, for this comparison, the equivalent Mask2Former [9] model is pre-trained with a significantly larger ImageNet-22k dataset, whereas our model is pre-trained on ImageNet-1k.
@@ IV-B2 ADE20K Dataset
We provide quantitative results on the ADE20K dataset in Table II. We choose one of the most commonly used crop sizes (640$\times$640) and provide results with both the ResNet-50 and ConvNeXt-L backbones.
In the small model comparison, we achieve the highest FPS and the second lowest # params (behind the surprisingly small RegProxy model), while outperforming RegProxy. In the large model comparison, our model is the smallest and fastest, and again outperforms RegProxy. Here, our model underperforms compared to the largest MaskFormer model, but achieves slightly better performance than the MaskFormer model at the equivalent latency budget. We also note that these larger MaskFormer models were also pre-trained with ImageNet-22k vs ours pre-trained with ImageNet-1k.
@@ IV-C2 Latency
As seen in Tables I and  II, our method provides a significant speed-up in latency compared to prior works. As the main processing of our model operates on the small superpixel space, this allows for a large reduction in model complexity, while even improving performance for the ResNet-50 backbone on Cityscapes case.

### v2

@@ Abstract
Semantic segmentation, which aims to classify every pixel in an image, is a key task in machine perception, with many applications across robotics and autonomous driving. Due to the high dimensionality of this task, most existing approaches use local operations, such as convolutions, to generate per-pixel features. However, these methods are typically unable to effectively leverage global context information due to the high computational costs of operating on a dense image. In this work, we propose a solution to this issue by leveraging the idea of superpixels, an over-segmentation of the image, and applying them with a modern transformer framework. In particular, our model learns to decompose the pixel space into a spatially low dimensional superpixel space via a series of local cross-attentions. We then apply multi-head self-attention to the superpixels to enrich the superpixel features with global context and then directly produce a class prediction for each superpixel. Finally, we directly project the superpixel class predictions back into the pixel space using the associations between the superpixels and the image pixel features. Reasoning in the superpixel space allows our method to be substantially more computationally efficient compared to convolution-based decoder methods. Yet, our method achieves state-of-the-art performance in semantic segmentation due to the rich superpixel features generated by the global self-attention mechanism. Our experiments on Cityscapes and ADE20K demonstrate that our method matches the state of the art in terms of accuracy, while outperforming in terms of model parameters and latency.
@@ I Introduction
Figure 1: Our method enables efficient segmentation of high-resolution camera images by learning to decompose the images into a set of superpixels. Specifically, we oversegment the image pixels into a small set of soft superpixels (left) via a series of local cross attentions. The superpixels are then refined via a set of multi-head self attentions, and directly classified. Finally, we fuse the class predictions with the superpixel-pixel associations to produce a dense semantic segmentation (right).
Figure 2: Our proposed Superpixel Transformer architecture. Given an image, we first generate hypercolumn features with an off-the-shelf encoder backbone. Our superpixel tokenization module uses a series of local dual-path cross-attentions to generate features for each superpixel. The superpixel features are then enriched by several multi-head self-attention (MHSA) layers to produce a class prediction for each superpixel, while the associations between each superpixel and pixel feature are computed from their respective features. Finally, the superpixel class predictions are unfolded into the dense pixel space using the associations. Note that the figure illustrates a hard assignment between pixels and superpixels for simplicity, while in practice we apply a differentiable soft assignment.
@@
Operating on the superpixel level provides a number of notable benefits. Conventional pixel-based approaches are limited by the high dimensionality of the pixel space, making global self-attention computationally intractable. Numerous approaches such as axial attention [42] or window attention [31] have been developed to work around these issues by relaxing the global attention to a local one. By over-segmenting the image into a small set of superpixels, we are able to efficiently apply global self-attention on the superpixels, providing full global context to the superpixel features, even when reasoning about high-resolution images. Despite applying global self-attention (vs conventional convolutional neural networks) in our model, our method is more efficient than existing methods due to the low dimensionality of the superpixel space. Finally, we directly produce semantic classes for the superpixel features, and then back-project the predicted classes onto the image space using the superpixel-pixel associations.
@@
  • A novel network architecture that uses local cross-attention to significantly reduce the spatial dimensionality of pixel features to a small set of superpixel features, enabling learning the global context between them and the direct classification of each superpixel.
  • A superpixel association and unfolding scheme that projects each superpixel class prediction back to a dense pixel segmentation, discarding the CNN pixel decoder.
@@ II Related Work
Transformers are also used as the decoders [2] for image segmentation. A popular design is to generate masks embedding vectors from object queries and then multiply them with the pixel features to generate masks [39, 44]. For example, MaX-DeepLab [41] proposes an end-to-end mask transformer framework that directly predicts class-labeled object masks. Segmenter [38] and MaskFormer [12] tackle semantic segmentation from the view of mask classification. K-Net [52] generates segmentation masks by a group of learnable kernels. Inspired by the similarity between mask transformers and clustering algorithms [33], clustering-based mask transformers are proposed to segment images [49, 48, 50]. Deformable transformer [57] is also used for improving the image segmentation as in Panoptic SegFormer [30] and Mask2Former [11].
Similar to this work, Region Proxy [53] (RegProxy) also incorporates the idea of superpixels into a deep segmentation network by using a CNN decoder to learn the association between each pixel and superpixel. However, RegProxy uses features on the regular pixel grid to represent each superpixel, and on which to apply self-attention. In comparison, we apply a set of learned weights, which correspond to each superpixel, and use cross-attention with the pixel features to compute the pixel-superpixel associations. Our experiments demonstrate that our methodology provides significant performance improvements.
In summary, all of the prior works that apply transformers to segmentation have, in some way, relied on a dense CNN decoder to generate the final dense features, and then combined these features with an attention mechanism to improve performance. Our method, in comparison, uses cross attention to reduce the image into a small set of superpixels, and only applies self-attention in this superpixel space in the decoder. This allows our method to operate on a significantly lower dimensional space (often $32^{2}\times$ smaller than the image resolution), while utilizing the benefits that come with global self-attention to achieve state-of-the-art performance.
@@ III Method
    Superpixel Classification: A series of multi-head self-attention layers to refine the superpixel features and produce a semantic class for each superpixel.
@@
    Superpixel Association: Associating the predicted superpixel classes and pixel features to obtain the final dense semantic segmentation.
@@ III-A Pixel Feature Extraction
Typical convolutional neural networks, such as ResNet [21] and ConvNeXt [32], are employed as the encoder backbone. On top of the encoder output, we apply a multi-layer perceptron (MLP), and bilinear resize to the features after stage-1 (stride 2), stage-3 (stride 8), and stage-5 (stride 32). The multi-scale features are combined with addition to form hypercolumn features [20]. Each pixel feature is represented by their corresponding hypercolumn features, which are fed to the following Superpixel Tokenization module.
@@ III-B Superpixel Tokenization
where $\mathcal{N}(x)$ denotes the neighborhood of $x$ and $q$, $k$ and $v$ are the query, key and value, generated applying a MLP to each respective feature plus an additive learned position embedding. For each superpixel neighborhood corresponding to a superpixel, we share the same set of position embeddings. For each superpixel $S_{i}$, there are $9\cdot h\cdot w$ pixel neighbors, where $[h,w]$ is the size of the patch covered by one superpixel, while each pixel has 9 superpixel neighbors. We illustrate this neighborhood in Figure 3. The local dual-path cross-attention repeats $n$ times to generate the output superpixel, $S_{i}^{t_{n}}$ and pixel, $I_{p}^{t_{n}}$, features.
This local dual-path cross-attention serves three purposes:
@@
III-C Superpixel Classification
[…]
@@ III-D Superpixel Association
To project the superpixel class predictions back into the pixel space, we use the outputs of the Superpixel Tokenization module, $I_{p}^{t_{n}}$ and $S_{i}^{t_{n}}$, to compute the association between each pixel and its 9 neighboring superpixels:
@@
The final dense semantic segmentation, $Y$, is then computed at each pixel, $p$, as the combination of each predicted superpixel class from the Superpixel Classification module, $C_{i}$, weighted by the above associations:
$Y_{p}=$ $\sum_{i\in\mathcal{N}(p)}Q_{pi}\cdot C_{i}.$ (6)
@@
MaskFormer [12] ResNet-50 [21] - - - 78.5
Mask2Former[11] ResNet-50 [21] - - - 79.4
Panoptic-DeepLab [10] ResNet-50 [21] 43M 517G - 78.7
$*{}^{*}$ [53] ViT-S [17] 23M 270G - 79.8
$k$${}^{\dagger}$ [50] ResNet-50 [21] 56M 434G 9.0 79.7
SP-Transformer ResNet-50 [21] 29M 253G 15.3 80.4
${}^{\ddagger}$ [11] Swin-L [31] - - - 83.3
$*{}^{*}$ [53] ViT-L/16 [17] 307M - - 81.4
SegFormer [47] MiT-B5 [47] 85M 1,448G 2.5 82.4
$k$${}^{\dagger}$ [50] ConvNeXt-L [32] 232M 1,673G 3.1 83.5
SP-Transformer ConvNeXt-L [32] 202M 1,557G 3.6 83.1
@@
Method Backbone Crop Params $\downarrow$ FLOPs $\downarrow$ FPS $\uparrow$ mIoU $\uparrow$
RegProxy [53] ViT-Ti/16 [17] 512 6M 3.9G 38.9 42.1
MaskFormer [12] ResNet-50 [21] 512 41M 53G 24.5 44.5
$k$${}^{\dagger}$ [50] ResNet-50 [21] 641 57M 75G 38.7 45.0
SP-Transformer ResNet-50 [21] 640 29M 78G 40.8 43.7
TABLE II: ADE20K val set results. We evaluate FLOPs and FPS with input $640\times 640$ for SP-Transformer on a Tesla V100-SXM2 GPU. Our method outperforms the prior RegProxy work, while remaining competitive with other prior works, and at the highest FPS. ${}^{\dagger}$kMaX-DeepLab is trained for panoptic segmentation.
@@ IV-A Implementation Details
Superpixel Classification 4 multi-head attention layers are applied in the superpixel classification stage, with 4 heads each, outputting 256 channels in each layer.
@@ IV-B1 Cityscapes Dataset
Table I compares the results of our Superpixel Transformer model to other transformer-based state-of-the-art models for semantic segmentation on the Cityscapes val set. In these experiments, we compare models with backbones roughly the same size as ResNet-50 and ConvNeXt-L. In addition to other semantic segmentation models, we also compare against the state of the art panoptic segmentation method, kMaX-DeepLab [50]. While this method is trained on a slightly different task, we find that, for computational cost, it is the most fair comparison, as our training schedule and pipeline is most similar to theirs.
With the smaller ResNet-50 backbone, our model is roughly half the size and latency of kMaX-DeepLab, while improving upon mIoU by 0.6 against the previous SOTA, RegProxy [53]. As the ResNet-50 backbone is relatively small, most existing models are dominated by the size of their decoders, which allows our model’s reduction in decoder size to have significant impacts on the overall size and performance, with a $70\%$ improvement in FPS compared to kMaX-DeepLab. In addition, we expect to see this effect even more for even smaller backbones.
For comparison, we also provided results with a larger ConvNeXt-L backbone. Here, our model has a similar absolute reduction in params and FLOPs, as compared to the equivalent kMaX-DeepLab model. However, as the model is largely dominated by the size of the backbone, the overall improvements are more modest. Nonetheless, our model is able to achieve near state-of-the-art performance, especially compared to the prior semantic segmentation methods, where we outperform most of the prior works, except for Mask2Former, where we are within 0.2mIoU. We note that, for this comparison, the equivalent Mask2Former [9] model is pre-trained with a significantly larger ImageNet-22k dataset, whereas our model is pre-trained on ImageNet-1k.
@@ IV-B2 ADE20K Dataset
We provide quantitative results on the ADE20K dataset in Table II. We choose one of the most commonly used crop sizes (640$\times$640) and provide results for the ResNet-50 backbone.
For ADE20K, we achieve the highest FPS and the second lowest # params (behind the surprisingly small RegProxy model), while outperforming RegProxy.
@@ IV-C2 Latency
As seen in Tables I and  II, our method provides a significant improvement in FPS. As the main processing of our model operates on the small superpixel space, this allows for a large reduction in model complexity, while achieving state of the art performance.